# Tensorization of neural networks for improved privacy and interpretability

### v2

@@ 4.2.2. Tensorization
from which we can define $G_{1}^{x_{1}}=\bra{0}G^{x_{1}}$ and $G_{n}^{x_{n}}=G^{x_{n}}\ket{0}$.
To apply TT-RSS, we fix the input dimension and rank as defined by the AKLT state, i.e., $d=3$ and $r=2$ (with $c=1-10^{-5}$). Since the function has a discrete domain, we use the trivial embedding for all sites. For pivot selection, we assume no knowledge of the state, and thus, we sample random configurations $(x_{1},\dots,x_{n})$ from the distribution defined by the AKLT state. While we assume no access to the TT representation of the state during the entire process, we leverage it for computational efficiency when obtaining the pivots, as it enables direct sampling. Alternatively, this step could be performed via Monte Carlo sampling, as required for NNQS. To analyze the performance of TT-RSS with different numbers of pivots, we use values of $N$ ranging from 2 to 14. Additionally, we run the experiments for three different values of $n$: 100, 200, and 500. For each configuration, TT-RSS is applied 10 times.
Upon tensorization, we obtain approximate versions of |$\ket{\Psi}$, given by
(4.10) |$\ket{\widehat{\Psi}}=\sum_{x_{1},\ldots,x_{n}}\widehat{G}_{1}^{x_{1}}\cdots\widehat{G}_{n}^{x_{n}}\ket{x_{1}\cdots x_{n}}.$
@@
where $W_{k}\in U(2)$ for all $k\in\{1,\ldots,n-1\}$.
To evaluate the accuracy of the reconstruction, we measure the fidelity between the recovered state |$\ket{\widehat{\Psi}}$ and the original state |$\ket{\Psi}$, as defined in Eq. 3.3. Remarkably, we observe that using as few as 12 pivots is sufficient to achieve a fidelity of 1, meaning that the AKLT state is recovered exactly, regardless of the system size $n$ (see Fig. 4.4).
@@ 4.2.3. Order parameter
(4.13) $u_{g}^{\otimes n}\ket{\Psi}=\ket{\Psi}.$
This implies, by virtue of the Fundamental Theorem of MPS [43], that there must exist a TI representation of $u_{g}^{\otimes n}\ket{\Psi}$ given by tensors $H$, such that
@@
where the physical transformation $u_{g}$ induces a transformation on the virtual degrees of freedom, represented by $V_{g}$. These transformations are not unique; transformations of the form $\widetilde{V}_{g}=\beta(g)V_{g}$, with $\beta(g)\in U(1)$, also satisfy Eq. 4.14. Consequently, they yield projective representations of the symmetry group, which can be classified into equivalence classes. These equivalence classes form a group isomorphic to the second cohomology group, $H^{2}(\mathbb{Z}_{2}\times\mathbb{Z}_{2},U(1))$, which labels the inequivalent representations.
@@
(4.15) $G^{x_{k}}\cdots G^{x_{k+L}}U_{g}(x_{k},\ldots,x_{k+L},y_{k},\ldots,y_{k+L})G^{y_{k}}\cdots G^{y_{k+L}}\approx(\Lambda V_{g})\otimes(\Lambda\bar{V}_{g}),$
where $\Lambda=\sum_{i}\lambda_{i}\ket{i}\bra{i}$ |, with $\lambda_{i}=\frac{1}{\sqrt{2}}$ for $i\in[2]$ being the Schmidt values. This approximation converges exponentially fast with the blocking size $L$. For the subscript notation, we use the Pauli matrix index $i$ to label the different group elements and their corresponding representations, as shown in Table 4.1, that is, $g_{i}\leftrightarrow u_{i}\leftrightarrow V_{i}\leftrightarrow\sigma_{i}$.
@@
(4.16) $\mathcal{E}_{L}(\Psi):=\bra{\Psi}(U_{3}\otimes U_{3}\otimes\mathbb{I})\mathbb{F}_{13}(U_{1}\otimes U_{1}\otimes\mathbb{I})\ket{\Psi},$
@@
In the description of the order parameter $\mathcal{E}_{L}$, we assume that the TT is TI. However, as pointed out earlier, the result from TT-RSS after canonicalization still retains a $U(2)$ gauge freedom. Nevertheless, when computing the order parameter while accounting for the unitary transformations $W_{k}$ acting on the virtual indices as shown in Eq. 4.12, these transformations cancel out, yielding the same result as in the TI case.
Figure 4.4. Fidelity between the AKLT state |$\ket{\Psi}$ and the approximated TT representation obtained via TT-RSS, |$\ket{\widehat{\Psi}}$, along with the order parameter $16\mathcal{E}_{10}$ to distinguish the Haldane phase from the trivial phase. The right figure shows the time taken to perform TT-RSS in seconds. For each value of $N$, the decomposition is performed 10 times, displaying the mean values with error bars at ±$\pm 0.5\sigma$, where $\sigma$ denotes the standard deviation. Three configurations are considered with different numbers of variables, $n=100$, $n=200$, and $n=500$, represented by different colors.
In our experiment, we measure $16\mathcal{E}_{10}(\Psi)\approx\pm 1$ to distinguish the trivial from the non-trivial phase. For the AKLT state, which is known to be in the non-trivial phase due to the anti-commutation relations of the Pauli matrices, the order parameter is $-1$. This is the value we must observe from the reconstructed TT representation of the state as well. As shown in Fig. 4.4, when the fidelity is 1—indicating an exact reconstruction of the AKLT state—the order parameter attains the value of $-1$. However, since our method does not rely on prior knowledge of the exact representation, it is applicable to cases where the achieved fidelity is less than 1. Notably, our results demonstrate that the order parameter reaches $-1$ with only 6 pivots in TT-RSS, even when fidelity remains below 1.
@@ 5.1. Initialization
In all cases, we set $d=2$ and $r=5$. For the embeddings, we consider two options: (i) the polynomial embedding $\phi(x)=[1,x]$, as proposed in Ref. [41], and (ii) the unit embedding $\phi(x)=[\cos(\frac{\pi}{2}x),\sin(\frac{\pi}{2}x)]$, introduced in Ref. [40] and commonly used for machine learning tasks. For the initializations, we explore the following settings:
@@ 5.2. Compression
In deep learning, where models are composed of intricate combinations of linear layers followed by non-linear activations, TNs have typically been employed to reduce the number of parameters in the inner linear layers. Among the available tensor decomposition techniques, the most common approach involves converting each matrix in a linear layer into a Tensor Train Matrix (TTM) [55, 58, 59, 60, 61, 62], also referred to as a Matrix Product Operator in physics literature. Other decomposition methods, such as the CP or Tucker decompositions, have also been successfully applied [56, 57]. This kind of tensorized models are commonly referred to as Tensor Neural Networks (TNNs).
@@
For example, in the base BERT model [5], some linear layers consist of matrices of size $768\times 3072$. Both the storage and computational complexities of multiplying a vector by such a matrix scale with its dimensions. Specifically, storing the matrix requires 2,359,296 floating-point numbers, and the multiplication involves 2,359,296 floating-point operations.
On the other hand, if the same matrix were represented as a TTM, whose cores are of the form $G_{k}(\alpha_{k-1},x_{k},\alpha_{k},y_{k})$, where $y_{k}$ represents the output index of each core and all ranks are set to $r=10$, we would need to store only 3,850 numbers. This represents a reduction of more than 99.8% in the number of parameters. However, contracting this TTM with an input vector (not in TT form) would require 2,810,880 operations, which exceeds the computational cost of using the original matrix. This higher complexity arises even with a relatively small rank $r=10$, which may already lead to significant approximation errors. The increased computational cost is primarily due to the need to recover the intermediate results as a full vector in order to apply non-linear activations element-wise, as these activations cannot be applied directly in TT form. The calculations assume the following decompositions for the input and output dimensions: $d^{in}=(1,2,2,2,2,2,2,2,2,3,1)$ and $d^{out}=(2,2,2,2,2,2,2,2,2,2,3)$.
Figure 5.2. Comparison of memory and time efficiency between different tensorized models and various choices of ranks $r$. The x-axis represents the ratio of parameters/operations in the tensorized models to those in the original models. The y-axis marks the accuracy difference, where negative values indicate accuracy drops in the tensorized models compared to the originals. Each color corresponds to a different NN model being tensorized. The legend indicates the number of hidden layers, approximate number of parameters, and test accuracy of each model. Dashed lines represent TT models obtained via TT-RSS, while dotted lines denote layer-wise tensorized TNNs. Points connected by these lines indicate, from left to right, increasing ranks: $r=2$, $r=5$, and $r=10$, with corresponding increases in the number of parameters and operations.
@@ 6. Conclusions
CRediT authorship contribution statement
José Ramón Pareja Monturiol: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Writing - original draft, Writing - review & editing, Visualization, Project administration. Alejandro Pozas-Kerstjens: Conceptualization, Validation, Formal analysis, Writing - review & editing, Supervision, Project administration. David Pérez-García: Conceptualization, Validation, Formal analysis, Writing - review & editing, Supervision, Project administration.
Declaration of interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Declaration of generative AI in the writing process
During the preparation of this work the authors used ChatGPT in order to improve the readability and language of the manuscript. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
@@ Acknowledgments
This work is supported by the Spanish Ministry of Science and Innovation MCIN/AEI/10.13039 /501100011033 (CEX2023-001347-S, CEX2019-000904-S, CEX2019-000904-S-20-4, PID2020-1135 23GB-I00, PID2023-146758NB-I00), Comunidad de Madrid (QUITEMAD-CM P2018/TCS-4342), Universidad Complutense de Madrid (FEI-EU-22-06), the CSIC Quantum Technologies Platform PTI-001, the Swiss National Science Foundation (grant number 224561), and the Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call – Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2026 Agenda. This research was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science, and Economic Development, and by the Province of Ontario through the Ministry of Colleges and Universities.

### v3

@@ 4.2.2. Tensorization
[…]
To apply TT-RSS, we fix the input dimension and rank as defined by the AKLT state, i.e., $d=3$ and $r=2$ (with $c=1-10^{-5}$). Since the function has a discrete domain, we use the trivial embedding for all sites. For pivot selection, we assume no knowledge of the state, and thus, we sample random configurations $(x_{1},\dots,x_{n})$ from the distribution defined by the AKLT state. While we assume no access to the TT representation of the state during the entire process, we leverage it for computational efficiency when obtaining the pivots, as it enables direct sampling. Alternatively, this step could be performed via Monte Carlo sampling, as required for NNQS. To analyze the performance of TT-RSS with different numbers of pivots, we use values of NN ranging from 2 to 14. Additionally, we run the experiments for three different values of $n$: 100, 200, and 500. For each configuration, TT-RSS is applied 10 times.
[…]
@@
where $W_{k}\in\mathrm{U}(2)$ for all $k\in[n-1]$.
[…]
@@ 4.2.3. Order parameter
[…]
@@
where the physical transformation $u_{g}$ induces a transformation on the virtual degrees of freedom, represented by $V_{g}$. These transformations are not unique; transformations of the form $\widetilde{V}_{g}=\beta(g)V_{g}$, with $\beta(g)\in\mathrm{U}(1)$, also satisfy Eq. 4.14. Consequently, they yield projective representations of the symmetry group, which can be classified into equivalence classes. These equivalence classes form a group isomorphic to the second cohomology group, $H^{2}(\mathbb{Z}_{2}\times\mathbb{Z}_{2},\mathrm{U}(1))$, which labels the inequivalent representations.
@@
(4.15) $G^{x_{k}}\cdots G^{x_{k+L}}U_{g}(x_{k},\ldots,x_{k+L},y_{k},\ldots,y_{k+L})\bar{G}^{y_{k}}\cdots\bar{G}^{y_{k+L}}\approx(\Lambda V_{g})\otimes(\Lambda\bar{V}_{g}),$
where $|\Lambda=\sum_{i}\lambda_{i}\ket{i}\bra{i}$, with $\lambda_{i}=\frac{1}{\sqrt{2}}$ for $i\in[2]$ being the Schmidt values. This approximation converges exponentially fast with the blocking size $L$. For the subscript notation, we use the Pauli matrix index $i$ to label the different group elements and their corresponding representations, as shown in Table 4.1, that is, $g_{i}\leftrightarrow u_{i}\leftrightarrow V_{i}\leftrightarrow\sigma_{i}$.
@@
[…]
@@
In the description of the order parameter $\mathcal{E}_{L}$, we assume that the TT is TI. However, as pointed out earlier, the result from TT-RSS after canonicalization still retains a $\mathrm{U}(2)$ gauge freedom. Nevertheless, when computing the order parameter while accounting for the unitary transformations $W_{k}$ acting on the virtual indices as shown in Eq. 4.12, these transformations cancel out, yielding the same result as in the TI case.
[…]
In our experiment, we measure $16\mathcal{E}_{10}(\Psi)\approx\pm 1$ to distinguish the trivial from the non-trivial phase. For the AKLT state, which is known to be in the non-trivial phase due to the anti-commutation relations of the Pauli matrices, the order parameter is 1-1. This is the value we must observe from the reconstructed TT representation of the state as well. As shown in Fig. 4.4, when the fidelity is 1—indicating an exact reconstruction of the AKLT state—the order parameter attains the value of 1-1. However, since our method does not rely on prior knowledge of the exact representation, it is applicable to cases where the achieved fidelity is less than 1. Notably, our results demonstrate that the order parameter reaches 1-1 with only 6 pivots in TT-RSS, even when fidelity remains below 1.
@@ 5.1. Initialization
For the embeddings, we consider two options: (i) the polynomial embedding $\phi(x)=[1,x]$, as proposed in Ref. [41], and (ii) the unit embedding $\phi(x)=[\cos(\frac{\pi}{2}x),\sin(\frac{\pi}{2}x)]$, introduced in Ref. [40] and commonly used for machine learning tasks. In all cases, we set $d=2$ and $r=5$. For the initializations, we explore the following settings:
@@ 5.2. Compression
In deep learning, where models are composed of intricate combinations of linear layers followed by non-linear activations, TNs have typically been employed to reduce the number of parameters in the inner linear layers. Among the available tensor decomposition techniques, the most common approach involves converting each matrix in a linear layer into a Matrix Product Operator, also referred to as Tensor Train Matrix (TTM) [55, 58, 59, 60, 61, 62]. Other decomposition methods, such as the CP or Tucker decompositions, have also been successfully applied [56, 57]. This kind of tensorized models are commonly referred to as Tensor Neural Networks (TNNs).
@@
For example, in the base BERT model [5], some linear layers consist of matrices of size $768\times 3\,072$. Both the storage and computational complexities of multiplying a vector by such a matrix scale with its dimensions. Specifically, storing the matrix requires 2 359 2962\,359\,296 floating-point numbers, and the multiplication involves 2 359 2962\,359\,296 floating-point operations.
On the other hand, if the same matrix were represented as a TTM, whose cores are of the form $G_{k}(\alpha_{k-1},x_{k},\alpha_{k},y_{k})$, where $y_{k}$ represents the output index of each core and all ranks are set to $r=10$, we would need to store only 3 8503\,850 numbers. This represents a reduction of more than 99.8% in the number of parameters. However, contracting this TTM with an input vector (not in TT form) would require 2 810 8802\,810\,880 operations, which exceeds the computational cost of using the original matrix. This higher complexity arises even with a relatively small rank $r=10$, which may already lead to significant approximation errors. The increased computational cost is primarily due to the need to recover the intermediate results as a full vector in order to apply non-linear activations element-wise, as these activations cannot be applied directly in TT form. The calculations assume the following decompositions for the input and output dimensions: $d^{in}=(1,2,2,2,2,2,2,2,2,3,1)$ and $d^{out}=(2,2,2,2,2,2,2,2,2,2,3)$.
@@
Figure 5.2. Comparison of memory and time efficiency between different tensorized models and various choices of ranks $r$. The x-axis represents the ratio of parameters/operations in the tensorized models to those in the original models. The y-axis marks the accuracy difference, where negative values indicate accuracy drops in the tensorized models compared to the originals. Each color corresponds to a different NN model being tensorized. The legend indicates the number of hidden layers, approximate number of parameters, and test accuracy of each model. Dashed lines represent TT models obtained via $T$-RSS, while dotted lines denote layer-wise tensorized TNNs. Points connected by these lines indicate, from left to right, increasing ranks: $r=2$, $r=5$, and $r=10$, with corresponding increases in the number of parameters and operations.
@@ Acknowledgments
This work is supported by the Spanish Ministry of Science and Innovation MCIN/AEI/10.13039/ 501100011033 (CEX2023-001347-S, CEX2019-000904-S, CEX2019-000904-S-20-4, PID2020-1135 23GB-I00, PID2023-146758NB-I00), Comunidad de Madrid (QUITEMAD-CM P2018/TCS-4342, TEC-2024/ COM-84-QUITEMAD-CM), Universidad Complutense de Madrid (FEI-EU-22-06), the CSIC Quantum Technologies Platform PTI-001, the Swiss National Science Foundation (grant number 224561), and the Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call – Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2026 Agenda. This research was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science, and Economic Development, and by the Province of Ontario through the Ministry of Colleges and Universities.